# FlowSpec: Continuous Pipelined Speculative Decoding for Efficient Distributed LLM Inference

### v1

@@ Abstract
Distributed inference serves as a promising approach to enabling the inference of large language models (LLMs) at the network edge. It distributes the inference process to multiple devices to ensure that the LLMs can fit into the device memory. Recent pipeline-based approaches have the potential to parallelize communication and computation, which helps reduce inference latency. However, the benefit diminishes when the inference request at the network edge is sparse, where pipeline is typically at low utilization. To enable efficient distributed LLM inference at the edge, we propose FlowSpec, a pipeline-parallel tree-based speculative decoding framework. FlowSpec incorporates three key mechanisms to improve decoding efficiency: 1) score-based step-wise verification prioritizes more important draft tokens to bring earlier accpeted tokens; 2) efficient draft management to prune invalid tokens while maintaining correct causal relationship during verification; 3) dynamic draft expansion strategies to supply high-quality speculative inputs. These techniques work in concert to enhance both pipeline utilization and speculative efficiency. We evaluate FlowSpec on a real-world testbed with other baselines. Experimental results demonstrate that our proposed framework significantly improves inference speed across diverse models and configurations, achieving speedup ratios 1.36$\times$-1.77$\times$ compared to baselines. Our code is publicly available at https://github.com/Leosang-lx/FlowSpec#
@@ 4.2 Main Results
Table 1: Main Results, which contain the performance comparison between FlowSpec and the baselines across 6 datasets under two sampling settings (Temperature = 0 or 1, 0 means greedy sampling). We select 20 samples for each dataset and limit the length of the generated sequences to 128. V and L2 are short for LLaMA2-Chat and Vicuna-v1.3, respectively. 7B and 13B denote the number of parameters of the respective models.
@@
V 13B Naive PP 0.98 1.00$\times$ 1.23 1.00$\times$ 1.20 1.00$\times$ 1.06 1.00$\times$ 0.62 1.00$\times$ 0.81 1.00$\times$ 0.98 1.00$\times$
PipeDec 1.43 1.46$\times$ 1.54 1.25$\times$ 1.54 1.28$\times$ 1.48 1.40$\times$ 1.17 1.89$\times$ 1.33 1.64$\times$ 1.42 1.49$\times$
FlowSpec 1.66 1.69$\times$ 2.05 1.67$\times$ 2.02 1.68$\times$ 1.80 1.70$\times$ 1.09 1.76$\times$ 1.45 1.74$\times$ 1.67 1.70$\times$
V 7B Naive PP 6.58 1.00$\times$ 5.29 1.00$\times$ 6.30 1.00$\times$ 5.95 1.00$\times$ 3.46 1.00$\times$ 4.07 1.00$\times$ 5.28 1.00$\times$
PipeDec 6.99 1.06$\times$ 6.71 1.27$\times$ 7.19 1.14$\times$ 6.95 1.17$\times$ 5.14 1.49$\times$ 5.88 1.44$\times$ 6.48 1.23$\times$
FlowSpec 8.93 1.36$\times$ 7.55 1.43$\times$ 8.60 1.37$\times$ 8.28 1.39$\times$ 4.74 1.37$\times$ 5.78 1.42$\times$ 7.31 1.38$\times$
L2 13B Naive PP 1.40 1.00$\times$ 1.60 1.00$\times$ 1.35 1.00$\times$ 1.22 1.00$\times$ 1.09 1.00$\times$ 1.11 1.00$\times$ 1.30 1.00$\times$
PipeDec 1.62 1.16$\times$ 1.64 1.02$\times$ 1.60 1.19$\times$ 1.56 1.28$\times$ 1.43 1.31$\times$ 1.52 1.37$\times$ 1.56 1.20$\times$
FlowSpec 2.29 1.64$\times$ 2.69 1.68$\times$ 2.26 1.67$\times$ 2.04 1.67$\times$ 1.81 1.66$\times$ 1.89 1.70$\times$ 2.16 1.66$\times$
L2 7B Naive PP 7.05 1.00$\times$ 7.56 1.00$\times$ 6.24 1.00$\times$ 6.13 1.00$\times$ 4.56 1.00$\times$ 5.28 1.00$\times$ 6.14 1.00$\times$
PipeDec 7.27 1.03$\times$ 7.36 0.97$\times$ 7.12 1.14$\times$ 7.08 1.15$\times$ 5.79 1.27$\times$ 6.67 1.26$\times$ 6.88 1.12$\times$
FlowSpec 9.43 1.34$\times$ 10.07 1.33$\times$ 8.71 1.40$\times$ 8.28 1.35$\times$ 6.10 1.34$\times$ 7.33 1.39$\times$ 8.32 1.35$\times$
@@
V 13B Naive PP 0.79 1.00$\times$ 0.73 1.00$\times$ 0.81 1.00$\times$ 0.88 1.00$\times$ 0.70 1.00$\times$ 0.70 1.00$\times$ 0.77 1.00$\times$
PipeDec 1.37 1.73$\times$ 1.26 1.73$\times$ 1.40 1.73$\times$ 1.21 1.38$\times$ 1.34 1.91$\times$ 1.32 1.89$\times$ 1.32 1.71$\times$
FlowSpec 1.38 1.75$\times$ 1.36 1.86$\times$ 1.36 1.68$\times$ 1.47 1.67$\times$ 1.17 1.67$\times$ 1.26 1.80$\times$ 1.33 1.73$\times$
V 7B Naive PP 4.12 1.00$\times$ 4.15 1.00$\times$ 3.71 1.00$\times$ 4.56 1.00$\times$ 3.13 1.00$\times$ 3.47 1.00$\times$ 3.86 1.00$\times$
PipeDec 6.19 1.50$\times$ 6.34 1.53$\times$ 5.92 1.60$\times$ 6.28 1.38$\times$ 5.12 1.64$\times$ 5.64 1.63$\times$ 5.92 1.53$\times$
FlowSpec 5.76 1.40$\times$ 5.77 1.39$\times$ 5.47 1.47$\times$ 6.17 1.35$\times$ 4.63 1.48$\times$ 5.16 1.49$\times$ 5.49 1.42$\times$
L2 13B Naive PP 1.36 1.00$\times$ 1.51 1.00$\times$ 1.22 1.00$\times$ 1.20 1.00$\times$ 1.02 1.00$\times$ 1.15 1.00$\times$ 1.24 1.00$\times$
PipeDec 1.61 1.18$\times$ 1.66 1.10$\times$ 1.60 1.31$\times$ 1.55 1.29$\times$ 1.45 1.42$\times$ 1.55 1.35$\times$ 1.57 1.27$\times$
FlowSpec 2.18 1.60$\times$ 2.29 1.52$\times$ 2.09 1.71$\times$ 1.99 1.66$\times$ 1.67 1.64$\times$ 1.82 1.58$\times$ 2.01 1.62$\times$
L2 7B Naive PP 6.46 1.00$\times$ 6.48 1.00$\times$ 5.95 1.00$\times$ 5.87 1.00$\times$ 4.17 1.00$\times$ 5.04 1.00$\times$ 5.66 1.00$\times$
PipeDec 7.11 1.10$\times$ 7.06 1.09$\times$ 7.25 1.22$\times$ 7.13 1.21$\times$ 5.78 1.39$\times$ 6.82 1.35$\times$ 6.86 1.21$\times$
FlowSpec 8.53 1.32$\times$ 8.63 1.33$\times$ 8.43 1.42$\times$ 7.54 1.28$\times$ 5.79 1.39$\times$ 7.00 1.39$\times$ 7.65 1.35$\times$
@@
Table 1 presents the results of FlowSpec and the baselines across multiple tasks on 4 LLM models. From the table, we observe that FlowSpec significantly outperforms both Naive PP and PipeDec on a variety of tasks. Across multiple tasks, FlowSpec achieves an average speedup of up to 1.73× compared to Naive PP.
@@
Temperature = 0. Under the greedy sampling strategy, FlowSpec’s average performance on each model surpasses that of PipeDec and Naive PP. Specifically, on the Vicuna-v1.3 13B model, it achieves a 1.70× speedup over Naive PP, whereas PipeDec only attains a 1.49× speedup under the same conditions. More specifically, FlowSpec achieves a 1.76× speedup on the instruction understanding (Alpaca) dataset with the Vicuna-v1.3 13B model, demonstrating the effectiveness of our method. It is evident that FlowSpec delivers nearly consistent speedups across different tasks. For example, on the LLaMA2-Chat 13B model it achieves an approximately 1.67× speedup, demonstrating that our method is effective and exhibits both robustness and stability under a variety of conditions. Overall, FlowSpec illustrates effectiveness and robustness at temperature = 0.
@@
Temperature = 1. Table 1 highlights FlowSpec’s remarkable advantage: on the Vicuna-v1.3 13B model, it runs 1.73× faster than Naive PP at temperature = 1. Under the same settings, it delivers a 1.86× boost on the code generation task (HumanEval). Moreover, it generates 7.65 tokens/s on the LLaMA2-Chat 7B on average, demonstrating outstanding efficacy. Owing to the uncertainty inherent in stochastic sampling, the acceleration achieved across different datasets and models exhibits appreciable variability. FlowSpec consistently delivers significant speedups across a variety of models and tasks.
@@
Under the Vicuna-v1.3 7B model as the base model with temperature=1, FlowSepc’s performance is marginally inferior to that of PipeDec. This is mainly due to the high decoding stochasticity and low draft acceptance rates from the draft model, leading to frequent cold starts in both approaches. Compared to FlowSpec, PipeDec’s layer-wise expansion strategy incurs slightly lower cold-start overhead, which may contribute to its modest performance gain in such cases. Nevertheless, FlowSpec stil demonstrates overall superiority and efficiency.
@@ 4.3 Ablation Study
Tree Pruning. The difference between Naive Pipeline Parallelism and Pruned Pipeline Parallelism highlights the effectiveness of tree pruning: enabling pruning yields a 1.25× speedup over Naive PP, demonstrating a clear performance improvement. Redundant computations in tree-based speculative decoding substantially degrade inference efficiency.

### v2

@@ Abstract
Distributed inference serves as a promising approach to enabling the inference of large language models (LLMs) at the network edge. It distributes the inference process to multiple devices to ensure that the LLMs can fit into the device memory. Recent pipeline-based approaches have the potential to parallelize communication and computation, which helps reduce inference latency. However, the benefit diminishes when the inference request at the network edge is sparse, where pipeline is typically at low utilization. To enable efficient distributed LLM inference at the edge, we propose FlowSpec, a pipeline-parallel tree-based speculative decoding framework. FlowSpec incorporates three key mechanisms to improve decoding efficiency: 1) score-based step-wise verification prioritizes more important draft tokens to bring earlier accpeted tokens; 2) efficient draft management to prune invalid tokens while maintaining correct causal relationship during verification; 3) dynamic draft expansion strategies to supply high-quality speculative inputs. These techniques work in concert to enhance both pipeline utilization and speculative efficiency. We evaluate FlowSpec on a real-world testbed with other baselines. Experimental results demonstrate that our proposed framework significantly improves inference speed across diverse models and configurations, achieving speedup ratios 1.28$\times$-1.79$\times$ compared to baselines. Our code is publicly available at https://github.com/Leosang-lx/FlowSpec#
@@ 4.2 Main Results
Table 1: Main Results, which contain the performance comparison between FlowSpec and the baselines across 6 datasets under two sampling settings (Temperature = 0 or 1, 0 means greedy sampling). We select 20 samples for each dataset and limit the length of the generated sequences to 256. V and L2 are short for LLaMA2-Chat and Vicuna-v1.3, respectively. 7B and 13B denote the number of parameters of the respective models.
@@
V 13B Naive PP 1.58 1.00$\times$ 1.56 1.00$\times$ 1.42 1.00$\times$ 1.25 1.00$\times$ 1.15 1.00$\times$ 1.03 1.00$\times$ 1.33 1.00$\times$
PipeDec 1.68 1.06$\times$ 1.70 1.09$\times$ 1.67 1.18$\times$ 1.62 1.30$\times$ 1.45 1.26$\times$ 1.51 1.47$\times$ 1.60 1.20$\times$
FlowSpec 2.55 1.61$\times$ 2.54 1.63$\times$ 2.35 1.65$\times$ 2.05 1.64$\times$ 1.82 1.58$\times$ 1.74 1.69$\times$ 2.18 1.64$\times$
V 7B Naive PP 7.69 1.00$\times$ 7.60 1.00$\times$ 6.83 1.00$\times$ 6.27 1.00$\times$ 4.81 1.00$\times$ 5.04 1.00$\times$ 6.37 1.00$\times$
PipeDec 7.61 0.99$\times$ 7.58 1.00$\times$ 7.49 1.10$\times$ 7.27 1.16$\times$ 6.19 1.29$\times$ 6.81 1.35$\times$ 7.16 1.12$\times$
FlowSpec 9.88 1.29$\times$ 9.47 1.28$\times$ 8.91 1.31$\times$ 8.28 1.32$\times$ 6.43 1.34$\times$ 7.17 1.42$\times$ 8.40 1.32$\times$
L2 13B Naive PP 1.45 1.00$\times$ 1.58 1.00$\times$ 1.37 1.00$\times$ 1.25 1.00$\times$ 1.12 1.00$\times$ 1.13 1.00$\times$ 1.31 1.00$\times$
PipeDec 1.64 1.13$\times$ 1.69 1.07$\times$ 1.63 1.19$\times$ 1.60 1.28$\times$ 1.47 1.29$\times$ 1.55 1.37$\times$ 1.59 1.21$\times$
FlowSpec 2.36 1.63$\times$ 2.52 1.60$\times$ 2.26 1.65$\times$ 2.10 1.68$\times$ 1.82 1.63$\times$ 1.92 1.70$\times$ 2.16 1.65$\times$
L2 7B Naive PP 6.86 1.00$\times$ 7.29 1.00$\times$ 6.23 1.00$\times$ 6.19 1.00$\times$ 4.77 1.00$\times$ 5.29 1.00$\times$ 6.10 1.00$\times$
PipeDec 7.33 1.07$\times$ 7.49 1.02$\times$ 7.25 1.16$\times$ 7.19 1.16$\times$ 6.17 1.29$\times$ 6.78 1.29$\times$ 7.03 1.15$\times$
FlowSpec 9.02 1.31$\times$ 9.51 1.30$\times$ 8.50 1.36$\times$ 8.49 1.37$\times$ 6.41 1.34$\times$ 7.46 1.41$\times$ 8.23 1.35$\times$
@@
V 13B Naive PP 1.31 1.00$\times$ 1.41 1.00$\times$ 1.17 1.00$\times$ 1.06 1.00$\times$ 1.01 1.00$\times$ 0.97 1.00$\times$ 1.16 1.00$\times$
PipeDec 1.64 1.25$\times$ 1.64 1.16$\times$ 1.59 1.36$\times$ 1.55 1.46$\times$ 1.41 1.40$\times$ 1.49 1.54$\times$ 1.55 1.34$\times$
FlowSpec 2.24 1.71$\times$ 2.36 1.67$\times$ 2.10 1.79$\times$ 1.85 1.75$\times$ 1.66 1.64$\times$ 1.57 1.62$\times$ 1.96 1.70$\times$
V 7B Naive PP 6.34 1.00$\times$ 6.31 1.00$\times$ 5.71 1.00$\times$ 5.19 1.00$\times$ 4.47 1.00$\times$ 4.38 1.00$\times$ 5.40 1.00$\times$
PipeDec 7.31 1.15$\times$ 7.38 1.17$\times$ 7.12 1.25$\times$ 6.87 1.32$\times$ 5.94 1.33$\times$ 6.56 1.50$\times$ 6.86 1.27$\times$
FlowSpec 8.26 1.30$\times$ 8.71 1.38$\times$ 7.88 1.38$\times$ 7.37 1.42$\times$ 5.96 1.33$\times$ 6.32 1.44$\times$ 7.42 1.37$\times$
L2 13B Naive PP 1.35 1.00$\times$ 1.52 1.00$\times$ 1.30 1.00$\times$ 1.20 1.00$\times$ 1.07 1.00$\times$ 1.15 1.00$\times$ 1.26 1.00$\times$
PipeDec 1.65 1.22$\times$ 1.70 1.12$\times$ 1.63 1.25$\times$ 1.59 1.32$\times$ 1.47 1.37$\times$ 1.58 1.37$\times$ 1.60 1.26$\times$
FlowSpec 2.32 1.72$\times$ 2.63 1.73$\times$ 2.22 1.71$\times$ 2.04 1.70$\times$ 1.72 1.61$\times$ 1.99 1.73$\times$ 2.15 1.71$\times$
L2 7B Naive PP 6.49 1.00$\times$ 6.64 1.00$\times$ 6.10 1.00$\times$ 5.90 1.00$\times$ 4.50 1.00$\times$ 5.24 1.00$\times$ 5.81 1.00$\times$
PipeDec 7.30 1.12$\times$ 7.17 1.07$\times$ 7.36 1.20$\times$ 7.22 1.22$\times$ 6.13 1.36$\times$ 6.95 1.32$\times$ 7.02 1.20$\times$
FlowSpec 8.53 1.31$\times$ 8.90 1.34$\times$ 8.37 1.37$\times$ 8.01 1.36$\times$ 6.14 1.36$\times$ 7.31 1.40$\times$ 7.88 1.36$\times$
@@
Table 1 presents the results of FlowSpec and the baselines across multiple tasks on 4 LLM models. From the table, we observe that FlowSpec significantly outperforms both Naive PP and PipeDec on a variety of tasks. Across multiple tasks, FlowSpec achieves an average speedup of up to 1.79$\times$ compared to Naive PP on the mathematical reasoning task.
@@
Temperature = 0. Under the greedy sampling strategy, FlowSpec’s average performance on each model surpasses that of PipeDec and Naive PP. Specifically, on the LLaMA2-Chat 13B model, it achieves a 1.65$\times$ speedup over Naive PP, whereas PipeDec only attains a 1.21$\times$ speedup under the same conditions. More specifically, FlowSpec achieves a 1.70$\times$ speedup on the question answering (Natural Questions) dataset with the LLaMA2-Chat 13B model as well, demonstrating the effectiveness of our method. It is evident that FlowSpec delivers nearly consistent speedups across different tasks. For example, on the Vicuna-v1.3 7B model, it approximately achieves the average speedup ratio of 1.32$\times$ across these tasks, demonstrating that our method is effective and exhibits both robustness and stability under a variety of conditions. Overall, FlowSpec illustrates effectiveness and robustness at temperature = 0.
@@
Temperature = 1. Table 1 highlights FlowSpec’s remarkable advantage: on the LLaMA2-Chat 13B model, it runs 1.71$\times$ faster than Naive PP at temperature = 1. On the Vicuna-v1.3 13B, it delivers a 1.79$\times$ boost on the math reasoning task (GSM8K). Moreover, it generates 7.88 tokens/s on the LLaMA2-Chat 7B on average, demonstrating outstanding efficacy. Owing to the uncertainty inherent in stochastic sampling, the acceleration achieved across different datasets and models exhibits appreciable variability.
@@
When executing text summarization on CNN/Daily Mail with temperature=1 on Vicuna-v1.3 7B, FlowSpec’s performance is marginally inferior to that of PipeDec. This is mainly due to the high decoding stochasticity and low draft acceptance rates from the draft model, leading to frequent cold starts in both approaches. Compared to FlowSpec, PipeDec’s layer-wise expansion strategy incurs slightly lower cold-start overhead, which may contribute to its modest performance gain in such cases. Nevertheless, FlowSpec demonstrates overall superiority and efficiency.
@@ 4.3 Ablation Study
Tree Pruning. The difference between Naive Pipeline Parallelism and Pruned Pipeline Parallelism highlights the effectiveness of tree pruning: enabling pruning yields a 1.25$\times$ speedup over Naive PP, demonstrating a clear performance improvement. Redundant computations in tree-based speculative decoding substantially degrade inference efficiency.